# On quasi-Albanese morphisms for log canonical Calabi-Yau pairs

### v1

@@ 3. Proof of the main theorems
The idea is from [Fuj24, Proof of Theorem 1.3, Step 2]. Let $A$ be the abelian part of $G$ and let $p:G\to A$ be the natural projection.
Case 1. $P$ is vertical over $A$. Then $P=p^{*}Q$ for some prime divisor $Q$ on $A$. Let $\mathbb{P}_{A}\to A$ be a $(\mathbb{P}_{1})^{d}$-bundle over $A$ such that $\mathbb{P}_{A}$ is a compactification of $G$, and $H=\mathbb{P}_{A}\setminus G$ is a snc divisor with $K_{\mathbb{P}_{A}}+H\sim 0$. Let $\overline{P}$ be the closure of $P$ in $\mathbb{P}_{A}$. Since any lc center of $(\mathbb{P}_{A},H)$ is horizontal over $A$, $\overline{P}$ contains no lc centers of $(\mathbb{P}_{A},H)$. Then $\overline{\kappa}(G\setminus P)=\kappa(K_{\mathbb{P}_{A}}+H+\overline{P})=\kappa(\overline{P})=\kappa(Q)>0$ by Lemma 17.
Case 2. $P$ is horizontal over $A$. Then there is a subgroup $\mathbb{C}^{*}\subset G$ such that $P$ dominates the quasi-abelian variety $G^{\prime}:=G/\mathbb{C}^{*}$. Let $F$ be a general fiber of $G\setminus P\to G^{\prime}$. Then $F$ is a curve with $\overline{\kappa}(F)>0$. Hence $\overline{\kappa}(G\setminus P)\geq\overline{\kappa}(F)+\overline{\kappa}(G^{\prime})=\overline{\kappa}(F)>0$ by Kawamata [Kaw78, Theorem 1] (see also Fujino [Fuj20, Chapter 5]).∎
Lemma 17.
@@
Theorem 2 is a special case of the following theorem.
Theorem 19.
Let $(X,D_{X})$ be a $\mathbb{Q}$-factorial dlt pair with $D_{X}$ reduced and $K_{X}+D_{X}\sim_{\mathbb{Q}}0$. Let $\alpha:X\setminus D\to G$ be the quasi-Albanese morphism. Then the conclusion in Theorem 2 holds.
@@
Keep the notations introduced in the proof of Theorem 3.
@@
(1) Let $P\subset G$ be a prime divisor and let $\overline{P}$ be its closure in $\mathbb{P}_{A}$. Then $\overline{P}\not\subset H$. Since $D_{W}^{v}=g^{-1}H$, there is a prime divisor $\overline{Q}_{W}\subset W$ such that $\overline{Q}_{W}\not\subset D_{W}$ and $\overline{Q}_{W}$ dominates $\overline{P}$. Let $\overline{Q}$ be the strict transform of $\overline{Q}_{W}$ on $X$. Then $\overline{Q}\not\subset D_{X}$ and $\overline{Q}$ dominates $P$.
(2) Let $Q\subset X\setminus D_{X}$ be a prime divisor and let $\overline{Q}$ be its closure in $X$. Then $\overline{Q}\not\subset D_{X}$ and its strict transform $\overline{Q}_{W}$ on $W$ has $\overline{Q}_{W}\not\subset D_{W}$, so $\overline{Q}_{W}\not\subset g^{-1}H$. Since $(W,D_{W})|_{g^{-1}G}\to G$ is locally stable and in particular equidimensional, we have $\operatorname{codim}\alpha(Q)=\operatorname{codim}g(\overline{Q}_{W})=1$.
@@
(3) Let $P\subset G$ be a prime divisor and let $\overline{P}$ be its closure in $\mathbb{P}_{A}$. Since $(W,D_{W})|_{g^{-1}G}\to G$ is locally stable, we have $(W,D_{W}+g^{*}\overline{P})$ is lc near the generic point of $\overline{P}$, so $(Z,D_{Z}-E_{Z}+f^{*}\overline{P})|_{f^{-1}G}$ is sub-lc near the generic point of $P$. By
$(K_{Z}+D_{Z}-E_{Z}+f^{*}\overline{P})|_{Z\setminus\pi^{-1}D_{X}}=(\pi|_{Z\setminus\pi^{-1}D_{X}})^{*}(K_{X\setminus D_{X}}+\alpha^{*}P)$
and $Z\setminus\pi^{-1}D_{X}\subset Z\setminus f^{-1}H=f^{-1}{G}$, we have $(X\setminus D_{X},\alpha^{*}P)$ is lc near the generic point of $P$. ∎
4. A remark on orbifold fundamental groups
This section is inspired by the work [GLM23]. Let $(X,D_{X}=\sum_{i}(1-\frac{1}{m_{i}})D_{i})$ be an lc pair with standard coefficients, where each $D_{i}$ is a prime divisor and $m_{i}\in\mathbb{Z}_{>0}\cup\infty$. For each $i$, let $\gamma_{i}$ be a small loop around $D_{i}$. The orbifold fundamental group $\pi_{1}^{\operatorname{orb}}(X,D_{X})$ is defined as the quotient of $\pi_{1}(X^{\operatorname{reg}}\setminus D_{X})$ by the normal subgroup generated by $\gamma_{i}^{m_{i}}$’s. Note that if $D_{X}$ is reduced, then $\pi_{1}^{\operatorname{orb}}(X,D_{X})=\pi_{1}(X^{\operatorname{reg}}\setminus D_{X})$.
Conjecture 20.
Let $(X,D_{X})$ be an lc pair with standard coefficients. If $K_{X}+D_{X}\sim_{\mathbb{Q}}0$, then $\pi_{1}^{\operatorname{orb}}(X,D)$ is virtually nilpotent.
Using the index-one cover and the Galois correspondence for orbifold fundamental groups (cf. [GLM23, Proposition A.11]), Conjecture 20 is reduced to the case $D_{X}$ is reduced. By the following Lemma, Conjecture 20 is further reduced to the $\mathbb{Q}$-factorial dlt case.
Lemma 21.
Let $(X,D_{X})$ and $(Y,D_{Y})$ be two projective lc pairs with $D_{X}$ and $D_{Y}$ reduced. If there is a birational contraction $p:X\dashrightarrow Y$ such that $D_{X}=p^{-1}_{*}D_{Y}+\operatorname{Ex}(p)$, then there is an isomorphism $\pi_{1}^{\operatorname{orb}}(X,D_{X})\cong\pi_{1}^{\operatorname{orb}}(Y,D_{Y})$.
@@ 4. A remark on orbifold fundamental groups
Let $Z\subset X$ and $Z^{\prime}\subset Y$ be closed subsets such that $X\setminus Z\cong Y\setminus Z^{\prime}$. Hence $X^{\operatorname{reg}}\setminus Z\cong Y^{\operatorname{reg}}\setminus Z^{\prime}$ and $X^{\operatorname{reg}}\setminus D_{X}\cup Z\cong Y^{\operatorname{reg}}\setminus D_{Y}\cup Z^{\prime}$. Since $p$ is a birational contraction, we have $\operatorname{codim}Z^{\prime}\geq 2$ and $D_{X}\cup Z=D_{X}\cup Z^{\prime\prime}$ for some $\operatorname{codim}\geq 2$ closed subset $Z^{\prime\prime}$. Hence $\pi_{1}(X^{\operatorname{reg}}\setminus D_{X})\cong\pi_{1}(X^{\operatorname{reg}}\setminus D_{X}\cup Z^{\prime\prime})\cong\pi_{1}(X^{\operatorname{reg}}\setminus D_{X}\cup Z)\cong\pi_{1}(Y^{\operatorname{reg}}\setminus D_{Y}\cup Z^{\prime})\cong\pi_{1}(Y^{\operatorname{reg}}\setminus D_{Y})$. ∎
Let $(X,D_{X})$ be a $\mathbb{Q}$-factorial dlt pair with $D_{X}$ reduced and $K_{X}+D_{X}\sim_{\mathbb{Q}}0$.
By a theorem due to Arapura and Nori [AN99, Corollary 4.10], to show that $\pi_{1}^{\text{orb}}(X,D_{X})=\pi_{1}(X^{\operatorname{reg}}\setminus D_{X})$ is virtually nilpotent, it suffices to show that it is polycyclic. Recall that a group $G$ is called polycyclic if there is a series of subgroups $1=G_{n}\triangleleft G_{n-1}\triangleleft\cdots\triangleleft G_{0}=G$ such that each $G_{i}/G_{i+1}$ is cyclic. The advantage of the class of polycyclic groups is that they are closed under passing to subgroups, quotients, and extensions (cf. [KM79, 4.4.3. Exercise]), which is crucial to inductive arguments.
Assume that $q(X,D_{X})>0$, and let $g:(W,D_{W})\to\mathbb{P}_{A}$ be the fibration given by Theorem 3. By Lemma 21, $(X,D_{X})$ and $(W,D_{W})$ have the same orbifold fundamental group. Let $W^{0}=g^{-1}G$ and let $F$ be a general fiber of $g$. Then $(F,{D_{W}}|_{F})$ is lc and $K_{F}+D_{W}|_{F}\sim_{\mathbb{Q}}0$. By the following exact sequence (cf. [GLM23, Lemma 2.18])
$\pi_{1}^{\operatorname{orb}}(F,D_{W}|_{F})\to\pi_{1}^{\operatorname{orb}}(W^{0},{D_{W}}|_{W^{0}})\to\pi_{1}(G)\to 1,$
we have $\pi_{1}^{\operatorname{orb}}(X,D_{X})\cong\pi_{1}^{\operatorname{orb}}(W,D_{W})\cong\pi_{1}^{\operatorname{orb}}(W^{0},{D_{W}}|_{W^{0}})$ is polycyclic if $\pi_{1}^{\operatorname{orb}}(F,D_{W}|_{F})$ is polycyclic. Hence, we have shown the following by induction on dimension.
Proposition 22.
To prove Conjecture 20, it suffices to prove $\mathbb{Q}$-factorial dlt pairs $(X,D_{X})$ with $D_{X}$ reduced and $q(X,D_{X})=0$ have polycyclic orbifold fundamental groups.

### v2

@@ 3. Proof of the main theorems
The idea is from [fujino2024quasi, Proof of Theorem 1.3, Step 2]. Let $A$ be the compact part of $G$ and let $p:G\to A$ be the natural projection.
Case 1. $P$ is vertical over $A$. Then $P=p^{*}Q$ for some prime divisor $Q$ on $A$. Let $(\mathbb{P}_{A},H)$ be the canonical compactification of $G$. Let $\overline{P}$ be the closure of $P$ in $\mathbb{P}_{A}$. Since any lc center of $(\mathbb{P}_{A},H)$ is horizontal over $A$, $\overline{P}$ contains no lc centers of $(\mathbb{P}_{A},H)$. Then $\overline{\kappa}(G\setminus P)=\kappa(K_{\mathbb{P}_{A}}+H+\overline{P})=\kappa(\overline{P})=\kappa(Q)>0$ by Lemma 3.4.
Case 2. $P$ is horizontal over $A$. Then there is a subgroup $\mathbb{C}^{*}\subset G$ such that $P$ dominates the quasi-abelian variety $G^{\prime}:=G/\mathbb{C}^{*}$. Let $F$ be a general fiber of $G\setminus P\to G^{\prime}$. Then $F$ is a curve with $\overline{\kappa}(F)>0$. Hence $\overline{\kappa}(G\setminus P)\geq\overline{\kappa}(F)+\overline{\kappa}(G^{\prime})=\overline{\kappa}(F)>0$ by Kawamata [kawamata1977addition, Theorem 1] (see also Fujino [fujino2020iitaka, Chapter 5]).∎
Lemma 3.4.
@@
Theorem 3.6 (A stronger dlt version of Theorem 1.4).
Let $(X,D_{X})$ be a projective $\mathbb{Q}$-factorial dlt pair with $K_{X}+D_{X}\sim_{\mathbb{Q}}0$. Let $\Delta_{X}\leq\lfloor D_{X}\rfloor$ be a reduced divisor. Let $\alpha:X\setminus\Delta_{X}\to G$ be the quasi-Albanese morphism. Let $A$ be the compact part of $G$. Let $(\mathbb{P}_{A},H)$ be the canonical compactification of $G$. Then there are
  • a projective $\mathbb{Q}$-factorial dlt pair $(W,D_{W})$ with $K_{W}+D_{W}\sim_{\mathbb{Q}}0$ and
such that the following holds.
  1. (1)
    The rational map $g:W\dashrightarrow\mathbb{P}_{A}$ is a fibration,
  2. (2)
    The canonical bundle formula for $(W,D_{W})\to\mathbb{P}_{A}$ writes $K_{W}+D_{W}\sim_{\mathbb{Q}}g^{*}(K_{\mathbb{P}_{A}}+H)$, and the restriction $(W,D_{W})|_{g^{-1}G}\to G$ is locally stable.
  3. (3)
    Let $\Delta_{W}=\mu^{-1}_{*}\Delta_{X}+\operatorname{Ex}(\mu)$. Then $\Delta^{v}_{W}=D_{W}^{v}=\lfloor D_{W}^{v}\rfloor=g^{-1}H$ and the morphism $W\setminus\Delta_{W}\to G$ is the quasi-Albanese morphism of $W\setminus\Delta_{W}$.
@@
Let $\pi:Z\to X$ be a log resolution such that the rational map $f:Z\dashrightarrow\mathbb{P}_{A}$ is a morphism. By Theorem 3.1(2), $f$ is a fibration. Let $D_{Z}=\pi^{-1}_{*}D_{X}+\operatorname{Ex}(\pi)$ and let $\Delta_{Z}=\pi^{-1}_{*}\Delta_{X}+\operatorname{Ex}(\pi)$. Then $K_{Z}+D_{Z}=\pi^{*}(K_{X}+D_{X})+E_{Z}$ for some effective and exceptional divisor $E_{Z}$ and $f^{-1}H\subset\Delta_{Z}\subset\lfloor D_{Z}\rfloor$.
By Theorem 3.2, there is a good minimal model $(W,D_{W})$ of $(Z,D_{Z})$ over $\mathbb{P}_{A}$ satisfying (1) and (2) and such that the birational map $\mu:(W,D_{W})\dashrightarrow(X,D_{X})$ is a crepant birational contraction with $D_{W}=\mu^{-1}_{*}D_{X}+\operatorname{Ex}(\mu)$.
By $\lfloor D_{W}^{v}\rfloor=g^{-1}H\subset\Delta_{W}^{v}\subset\lfloor D_{W}^{v}\rfloor$, we have $\Delta_{W}^{v}=\lfloor D_{W}^{v}\rfloor$. Let $W\setminus\Delta_{W}\to G^{\prime}$ be the quasi-Albanese morphism. Since $W\setminus\Delta_{W}\to G$ is a fibration, the induced morphism $G^{\prime}\to G$ is a fibration. By Lemmas 2.3 and 2.7, we have $\dim G^{\prime}=q(W,\Delta_{W})=q(X,\Delta_{X})=\dim G$. Hence the fibration $G^{\prime}\to G$ is an isomorphism, i.e., $W\setminus\Delta_{W}\to G$ is the quasi-Albanese morphism. ∎
Proof of Theorem 1.4.
Let $\pi:Z\to X$ be a log resolution such that $\pi^{-1}\operatorname{nklt}$ is a divisor and the rational map $f:Z\dashrightarrow\mathbb{P}_{A}$ is a morphism. By Theorem 3.1(3), $f$ is a fibration. Let $D_{Z}=\pi^{-1}_{*}D_{X}+\operatorname{Ex}(\pi)$. Then $K_{Z}+D_{Z}=\pi^{*}(K_{X}+D_{X})+E_{Z}$ for some effective and exceptional divisor $E_{Z}$ and $f^{-1}H\subset\pi^{-1}\operatorname{nklt}\subset\lfloor D_{Z}\rfloor$.
By Theorem 3.2, there is a good minimal model $(W,D_{W})$ of $(Z,D_{Z})$ over $\mathbb{P}_{A}$ satisfying (1) and (2) and such that the birational map $\mu:(W,D_{W})\dashrightarrow(X,D_{X})$ is a crepant birational contraction with $D_{W}=\mu^{-1}_{*}D_{X}+\operatorname{Ex}(\mu)$.
By the same reasoning as in the previous Theorem 3.6, we have $W\setminus\lfloor D_{W}\rfloor\to G$ is the quasi-Albanese morphism. ∎
Proof of Theorem 1.3.
Keep the notations introduced in the proof of Theorem 1.4.
@@
(1) Let $P\subset G$ be a prime divisor and let $\overline{P}$ be its closure in $\mathbb{P}_{A}$. Then $\overline{P}\not\subset H$. Since $\lfloor D_{W}^{v}\rfloor=g^{-1}H$, there is a prime divisor $\overline{Q}_{W}\subset W$ such that $\overline{Q}_{W}\not\subset\lfloor D_{W}\rfloor$ and $\overline{Q}_{W}$ dominates $\overline{P}$. Since $\operatorname{Ex}(\mu)\leq\lfloor D_{W}\rfloor$, $\overline{Q}:=\mu_{*}\overline{Q}_{W}$ is a prime divisor on $X$ such that $\overline{Q}\not\subset\lfloor D_{X}\rfloor$ (in particular $\overline{Q}\not\subset\operatorname{nklt}(X,D_{X})$) and $\overline{Q}$ dominates $P$.
@@
(2) Let $\Sigma\subset G$ be a reduced snc divisor and let $\overline{\Sigma}$ be its closure in $\mathbb{P}_{A}$. Since $(W,D_{W})|_{g^{-1}G}\to G$ is locally stable, we have $(W,D_{W}+g^{*}\overline{\Sigma})|_{g^{-1}G}$ is lc. Hence $(Z,D_{Z}-E_{Z}+f^{*}\overline{\Sigma})|_{f^{-1}G}$ is sub-lc. By
$(K_{Z}+D_{Z}-E_{Z}+f^{*}\overline{\Sigma})|_{Z\setminus\pi^{-1}\operatorname{nklt}}=(\pi|_{Z\setminus\pi^{-1}\operatorname{nklt}})^{*}(K_{X\setminus\operatorname{nklt}}+D_{X}|_{X\setminus\operatorname{nklt}}+\alpha^{*}\Sigma)$
and $Z\setminus\pi^{-1}\operatorname{nklt}\subset Z\setminus f^{-1}H=f^{-1}{G}$, we have $(X\setminus\operatorname{nklt}+D_{X}|_{X\setminus\operatorname{nklt}},\alpha^{*}\Sigma)$ is lc. Hence $\alpha:(X\setminus\operatorname{nklt},D_{X}|_{X\setminus\operatorname{nklt}})\to G$ is locally stable by [Kollar-Families, Corollary 4.55]. It is flat with slc fibers by [Kollar-Families, Definition-Theorem 4.7]
3.3. Proofs of Lemma 1.5 and Theorem 1.6
Lemma 3.7.
Let $(X,D_{X})$ be a projective log smooth pair with $D_{X}$ reduced. Write $D_{X}=\sum_{i=1}^{m}D_{i}$ as a sum of prime divisors. Then
$q(X,D_{X})-q(X)=\{(a_{1},\cdots,a_{m})\in\mathbb{C}^{m}\mid\sum_{i}a_{i}c_{1}(O_{X}(D_{i}))=0\text{ in }H^{1}(\Omega_{X}^{1})\},\text{ and}$
[…]
@@
Consider the exact sequence
$0\to\Omega_{X}^{1}\to\Omega_{X}^{1}(\log D_{X})\to\bigoplus_{i=1}^{m}O_{D_{i}}\to 0,$
and the associated long exact sequence:
$0\to H^{0}(\Omega_{X}^{1})\to H^{0}(\Omega_{X}^{1}(\log D_{X}))\to\bigoplus_{i=1}^{m}H^{0}(O_{D_{i}})\xrightarrow{\delta}H^{1}(\Omega_{X}^{1}).$
Since $\delta(1|_{D_{i}})=c_{1}(O_{X}(D_{i}))\in H^{1}(\Omega_{X}^{1})$ (cf. [HartshorneGTM52, Chapter III Exercise 7.4]), where $1|_{D_{i}}$ is the constant function on $D_{i}$ with value $1$, the Lemma follows. ∎
Lemma 3.8 (cf. [enwright2024log, Lemma 2.41]).
Let $(X,D_{X})$ be a projective lc pair. Let $\pi:Z\to X$ be a log resolution and let $D_{Z}=\pi^{-1}_{*}D_{X}+\operatorname{Ex}(\pi)$. Then $c(Z,D_{Z})=c(X,D_{X})$.
Proof.
Write $\operatorname{Ex}(\pi)=\sum_{i=1}^{m}E_{i}$ as a sum of prime divisors. Since $|\lfloor D_{Z}\rfloor|=|\lfloor D_{X}\rfloor|+m$, to prove the Lemma, it suffices to show that $\operatorname{rank}\langle\lfloor D_{Z}\rfloor\rangle=\operatorname{rank}\langle\lfloor D_{X}\rfloor\rangle+m$. Since $\pi_{*}:\langle\lfloor D_{Z}\rfloor\rangle\otimes\mathbb{Q}\to\langle\lfloor D_{X}\rfloor\rangle\otimes\mathbb{Q}$ is surjective and $E_{i}$’s form a basis of its kernel (by the negativity Lemma [KM98, 3.39]), the Lemma follows. ∎
Proof of Lemma 1.5.
Let $\pi:Z\to X$ be a log resolution such that $\pi^{-1}\operatorname{nklt}(X,D_{X})$ is a divisor. Let $D_{Z}=\pi^{-1}_{*}D_{X}+\operatorname{Ex}(\pi)$. By the previous two Lemmas, we have $c(X,D_{X})=c(Z,D_{Z})=\dim Z-(q(Z,D_{Z})-q(Z))$. By Lemma 2.7, $X\setminus\operatorname{nklt}(X,D_{X})$ and $Z\setminus\lfloor D_{Z}\rfloor$ have the isomorphic quasi-Albanese variety $G$. By Iitaka’s construction, we have $\dim G=q(Z,D_{Z})$ and $\dim A=q(Z)$. Hence $c(X,D_{X})=\dim X-(\dim G-\dim A)$. ∎
Proof of Theorem 1.6.
Let $\Delta_{X}\leq\lfloor D_{X}\rfloor$ be a reduced divisor. Let $\pi:Z\to X$ be a Hacon’s $\mathbb{Q}$-factorial dlt modification [Hacon2014ACC, Proposition 3.3.1]. Let $D_{Z}=\pi^{-1}_{*}D_{X}+\operatorname{Ex}(\pi)$ and let $\Delta_{Z}=\pi^{-1}_{*}\Delta_{X}+\operatorname{Ex}(\pi)$. Then $K_{Z}+D_{Z}=\pi^{*}(K_{X}+D_{X})$, $\kappa(K_{Z}+D_{Z})=0$ and $\Delta_{Z}\leq\lfloor D_{Z}\rfloor$.
By Lemma 3.8, we have $c(X,\Delta_{X})=c(Z,\Delta_{Z})$. By Proposition 1.5, we have $c(Z,\Delta_{Z})=\dim Z-(\dim G-\dim A)$, where $G$ and $A$ are the quasi-Albanese variety and the Albanese variety of $Z\setminus\Delta_{Z}$ and $Z$ respectively. By 3.1(2), we have $\dim Z\geq\dim G$. Hence $c(X,\Delta_{X})\geq 0$.
Assume that $c(X,\Delta_{X})=0$. Then $\dim Z=\dim G$ and $\dim A=0$. Hence $G\cong(\mathbb{C}^{*})^{n}$. The quasi-Albanese morphism $Z\setminus\Delta_{Z}\to(\mathbb{C}^{*})^{n}$ is birational, so $X$ is rational. Assume further that $K_{X}+D_{X}\sim_{\mathbb{Q}}0$. Then $K_{Z}+D_{Z}\sim_{\mathbb{Q}}0$. Set $((\mathbb{P}^{1})^{n},H):=(\mathbb{P}^{1}_{z_{1}}\times\cdots\times\mathbb{P}^{1}_{z_{n}},\sum_{i}((z_{i}=0)+(z_{i}=\infty)))$. By Theorem 3.6, there are
  • a $\mathbb{Q}$-factorial dlt pair $(W,D_{W})$ with $K_{W}+D_{W}\sim_{\mathbb{Q}}0$,
  • a crepant birational contraction $\mu:(W,D_{W})\dashrightarrow(Z,D_{Z})$ with $D_{W}=\mu^{-1}_{*}D_{Z}+\operatorname{Ex}(\mu)$, and
  • a birational morphism $g:(W,D_{W})\to((\mathbb{P}^{1})^{n},H)$ with $K_{W}+D_{W}\sim_{\mathbb{Q}}g^{*}(K_{(\mathbb{P}^{1})^{n}}+H)$, and $\Delta_{W}=\lfloor D_{W}\rfloor=D_{W}=g^{-1}H$, where $\Delta_{W}=\mu^{-1}_{*}\Delta_{Z}+\operatorname{Ex}(\mu)$.
Hence we have $\Delta_{Z}=\lfloor D_{Z}\rfloor=D_{Z}$, $\Delta_{X}=\lfloor D_{X}\rfloor=D_{X}$ and the four pairs $(X,D_{X})$, $(Z,D_{Z})$, $(W,D_{W})$ $((\mathbb{P}^{1})^{n},H)$ are crepant birational to each other.
Let $P\subset W$ be a prime divisor which is exceptional over $(\mathbb{P}^{1})^{n}$, then $P\subset g^{-1}H=D_{W}$. By $1=a(P,W,D_{W})=a(P,(\mathbb{P}^{1})^{n},H)$, we have $P$ is toric. By [Brown2018GeomToric, Lemma 2.3.2], we have $(W,D_{W})$, $(Z,D_{Z})$ and $(X,D_{X})$ are toric.
By taking a log resolution of $(Z,D_{Z})$ and Lemma 2.7, we known that $Z\setminus\lfloor D_{Z}\rfloor$ and $X\setminus\operatorname{nklt}(X,D_{X})$ have the same quasi-Albanese variety $(\mathbb{C}^{*})^{n}$. By Theorem 1.3, we have the quasi-Albanese morphism $X\setminus\operatorname{nklt}(X,D_{X})\to(\mathbb{C}^{*})^{n}$ is birational and isomorphic in codimension one. ∎